OpenStudy (rebeleco518):

Did a pre-test and got some wrong-help me understand? questions posted below-->

OpenStudy (rebeleco518):

Question 1 The map shows the location of a mall, library, and school in a city: Britney traveled from the school to the mall and then from the mall to the library. Alice traveled directly from the school to the library. How many more miles did Britney travel than Alice?  8 miles  9 miles  10 miles  12 miles-wrong

OpenStudy (rebeleco518):

Question 2 What is the volume of a sphere with a radius of 6 inches?  112π cubic inches  288π cubic inches  290π cubic inches-wrong  300π cubic inches

OpenStudy (rebeleco518):

Question 3 The picture below shows a pole and its shadow: What is the height of the pole?  112 centimeters  100 centimeters  98 centimeters  64 centimeters-wrong

OpenStudy (rebeleco518):

Question 4 Samuel filled the glasses shown below completely with water. The total amount of water that Samuel poured into the glasses is 60 cubic centimeters. What is the height of glass 1? Round your answer to the nearest tenth. (Use π = 3.14.) Note that all measurements are in centimeters and images are not drawn to scale.  0.2 centimeter  1.7 centimeters  3.9 centimeters  5.6 centimeters-wrong

OpenStudy (rebeleco518):

Question 5 The length of the hypotenuse of a right triangle is 137 units. The length of one leg of the triangle is 105. Gina wrote the following step to find the length of the unknown leg: Length of the unknown leg = 137^2 − 105^2 = 18,769 − 11,025 = 7,744 units Which statement best explains whether Gina's step is correct or incorrect?  It is incorrect because the length of the unknown side is the square root of 29,794.  It is incorrect because the length of the unknown side is the square root of 7,744.  It is correct because the length of the unknown side is the difference of the lengths of the sides.-wrong  It is correct because the length of the unknown side is the difference of the squares of the sides.

OpenStudy (rebeleco518):

Question 6 The length of the hypotenuse of a right triangle is 24 inches. If the length of one leg is 8 inches, what is the approximate length of the other leg?   16.0 inches -wrong  18.4 inches  22.6 inches   25.3 inches

OpenStudy (wmj259):

Easy stuff. Seems like you got all of them wrong?

OpenStudy (rebeleco518):

@jim_thompson5910 @hartnn @Kamille @phi

OpenStudy (rebeleco518):

yes i did. and the subject overall is easy, i agree, but i need to figure out what i did wrong.

OpenStudy (rebeleco518):

my work for #1 was this- 5^2+12^2=169 13^2=169 B=17 A=13 so i guessed on the question

OpenStudy (wmj259):

Actually for the distance you just add how far they moved. So Britney moved 5 north and 7 west. So 5+7=12. Alice moves from the school directly to the library. So that's when you use the Pythagorean theorem b/c you don't know the hypotenuse.

OpenStudy (wmj259):

Okay?

OpenStudy (rebeleco518):

how did you get 7 for going west?

OpenStudy (wmj259):

actually not 7, its 12 so 5+12=17

OpenStudy (wmj259):

my mistake.

OpenStudy (rebeleco518):

ok, so i was right about that part

OpenStudy (rebeleco518):

btw- b=brintney and A=alice

OpenStudy (wmj259):

Are you missing an option?

OpenStudy (rebeleco518):

nope

OpenStudy (wmj259):

Britney is 17 and Alice is 13.

OpenStudy (wmj259):

WOOWW, The graph counts by 2's

OpenStudy (wmj259):

Not 1's.

OpenStudy (rebeleco518):

so it would be 10! (right?)

OpenStudy (rebeleco518):

@wmj259 ?

OpenStudy (rebeleco518):

So I went back through and did a double take and here is what I got so far--> 1-10 2-idek 3-112 4-idek 5-b 6-22.5 Am I right/can I have help with 2 & 4?

OpenStudy (wmj259):

1 should be 8 miles.

OpenStudy (wmj259):

2.) \[V=\frac{ 4 }{ 3 }*\pi*r ^{3}\]

OpenStudy (wmj259):

3.) Pythagorean Theorem, solve for leg.